OpenStudy (anonymous):

in an electric circuit, the current passing through a conductor varies inversely with the resistance. Suppose that when the current is 5A (amperes), the resistance is 44 ohms. What is the resistance when the current is 2A

OpenStudy (anonymous):

Current is 5A Resistance is 44 ohms (sorry)

OpenStudy (sheraz12345):

Find voltage from first set of data and then use it to calculate resistance afterwards

OpenStudy (mathstudent55):

Direct variation: \(y = kx\) Inverse variation: \(y = \dfrac{k}{x} \) Use the given info to find k. Then write \(y = \dfrac{k}{x} \) with the correct value of k you found. Then enter x and find y.

OpenStudy (sheraz12345):

Ahh yes... That is proper mathematical modeling way (y) @mathstudent55

OpenStudy (anonymous):

@mathstudent55 what do I substitute in for y and k?

OpenStudy (aum):

Current and resistance are inversely related. Therefore, current = k / resistance ---- (1) We are given that when resistance = 44 ohms, current = 5 A Substitute in (1): 5 = k / 44 k = 5 * 44 = 220 Put k = 220 in (1) current = 220 / resistance ---- (2) What is the resistance when the current is 2A? Put current = 2 in equation (2) and find resistance: 2 = 220 / resistance 2 * resistance = 220 resistance = 220 / 2 = 110 ohms.

OpenStudy (aum):

A quicker way to do this problem follows: Since current and resistance are inversely related it implies that as current decreases from 5A to 2A, which is a factor of 2/5, the resistance will increase by a factor of 5/2 (because of inverse relation). Therefore, new resistance = 5/2 * old resistance = 5/2 * 44 = 5 * 22 = 110 ohms.

OpenStudy (anonymous):

Thank you all!

OpenStudy (aum):

You are welcome.